Here's what the problem is...
Given:AL=SK
Prove:AS=LK
Here are my statements leading to my proff:
1.AL=SK
2.LS-LS
3.AL+LS=SK+LS
4.AL+LS=AS
5.SK+LS=LK
6.AS=LK

Found 2 solutions by richwmiller, JBarnum:
Answer by richwmiller(9132)

It would be nice to know what this figure is and where these lines are that you speak of

Answer by JBarnum(2044)

Given:AL=SK
Prove:AS=LK
lets make these numbers so you can see it a bit better
{ex: A=1 L=8 S=2 K=4 so (1)(8)= (2)(4)}
1.AL=SK
(1x8)=(2x4) this is equal
2.LS-LS
(8x2)-(8x2) ?
3.AL+LS=SK+LS
(1x8)+(8x2)=(2x4)+(8x2) this is equal
4.AL+LS=AS
(1x8)+(8x2)=(1x2) => 8+16=2 => 24=2 this cant happen
5.SK+LS=LK
(2x4)+(8x2)=(8x4) => 8+16=32 => 24=32 this cant happen
6.AS=LK
(1x2)=(8x4) 2=32 cant happen
maybe im misunderstanding something
this is how it normally goes {ex: A=1 L=8 S=2 K=4}
AL=SK (1x8)=(2x4)
A=SK/L 1=(2X4)/8, A/S=K/L 1/2=4/8
L=SK/A 8=(2X4)/1, L/S=K/A 8/2=4/1
S=AL/K 2=(1X8)/4, S/A=L/K 2/1=8/4
K=AL/S 4=(1X8)/2, K/A=L/S 4/1=8/2
THESE ARE ALL EQUAL AND POSSIBLE
BUT IF A=K AND L=S ITS POSSIBLE
1=1 AND 2=2 1x2=1x2
AL=SK
A=K
L=S
AS=LK
HOPE SOME OF THIS HELPS